OpenStudy (anonymous):

Is there anyone here that can give me detailed help I have some questions that I just can't seem to get right.

OpenStudy (anonymous):

whats the q

OpenStudy (anonymous):

Find the vertex, line of symmetry, and the minimum or maximum value. f(x)=8-x^2

OpenStudy (amistre64):

yeah, specific questions help out

OpenStudy (anonymous):

I really just need the vertex.

OpenStudy (anonymous):

yea if I could get anyone one here to answer my questions.

OpenStudy (amistre64):

it straddles the y axis, 8 above origin

OpenStudy (amistre64):

-x^2 is a turned over bowl so its max is its vertex y part

OpenStudy (amistre64):

and the symmetry is the y axis since it never leaves its home

OpenStudy (anonymous):

soup spilled out

OpenStudy (amistre64):

3 beans soup all over my lap; just great

OpenStudy (anonymous):

so the coordinates for the vertex are (2,8) ?

OpenStudy (anonymous):

vertex is \[-\frac{b}{2a}\] for the first coordinate. in this case you might want to write \[f(x)=-x^2+8\] so b = 0 and the vertex is (0,8)

OpenStudy (anonymous):

Ah!

OpenStudy (amistre64):



OpenStudy (amistre64):

i keep forgettin we got a draw lol

OpenStudy (anonymous):

LOL! THANK YOU!!!!

OpenStudy (anonymous):